OpenStudy (anonymous):

Couple questions for Algebra 1 and up!!

OpenStudy (anonymous):

A couple of questions i like the sound of that hahahah

OpenStudy (anonymous):

\[6n^{-3}y/2n^{-1}y^{-3}\] haha yeah ive done these before but icant remember exactly...

OpenStudy (anonymous):

oh, im suppoed to simplify that

OpenStudy (anonymous):

oh gawd i hate those

OpenStudy (anonymous):

To start - write it as a fraction like it should be. Then realize negative exponents mean it's 1 over the exponent.

OpenStudy (amistre64):

just subtract your bottom exponents from your top exponents; then sort the madness

OpenStudy (anonymous):

I did write it as a fraction, i couldnt find the button for that now. Like i said, my teacher was out today and over the year we did the charpters out of order... like 123456789 12 11 10, and my teacher was out so a sub gave us the wrong thing or something

OpenStudy (anonymous):

get the reciprocal to remove the negatives because no polynomial can have a negative exponent. then simplify.

OpenStudy (anonymous):

amistre i have no clue what u are talking about :/

OpenStudy (amistre64):

the button is: type in, frac{}{} into the eq editor

OpenStudy (amistre64):

subtraction; its the opposite of addition :/

OpenStudy (amistre64):

\[\frac{n^{top}}{n^{bottom}}=n^{(top-bottom)}\]

OpenStudy (anonymous):

\[\frac{6n^{-3}y}{2n^{-1}y{-3}}\]

OpenStudy (anonymous):

thats y^{-3}

OpenStudy (amistre64):

so, the number part is just a normal fraction 6/2 the rest is n^(-3--1) y^(1--3)

OpenStudy (anonymous):

i know subtraction is the opposite of addition!!! i didnt know if you took away the bottom N when it simplified 6n/2n.:( i went to kindergarten :(

OpenStudy (amistre64):

:) just subtract the bottom stuff from the top stuff (exponentwise); after you simplfy see if you still got negative exponents; if so, drop them under the bar

OpenStudy (anonymous):

Let's start here: \[\frac{6n^{−3}y}{2n^{−1}y^{−3}}\] Personally I like writing negative exponents as a reciprocal first. The effect is moving it from bottom to top or top to bottom: \[\frac{6yn^{1}y^{3}}{2n^{3}}\] Then simplify: \[\frac{6n^{1}y^{4}}{2n^{3}}\] And subtract exponents as they are being divided \[\frac{3y^{4}}{n^{2}}\]

OpenStudy (anonymous):

Thank you! and i have another one... what is the degree of the polynomial mean?

OpenStudy (amistre64):

\[\frac{6n^{−3}y}{2n^{−1}y^{−3}}=3n^{(-3--1)}y^{(1--3)}\to\ 3n^{-2}y^4=\frac{3y^4}{n^2}{}\]

OpenStudy (amistre64):

the degree of a poly is the highest exponential count of a single term

OpenStudy (amistre64):

4xyz = degree 3 x^2 y^5 = degree 7 etc ..

OpenStudy (amistre64):

but only variables count for some reason

OpenStudy (anonymous):

And of course, terms are separated by + or -: \(4x^2y^5+4j^6-3x^2+5\) Can you find the degree of this polynomial?

OpenStudy (anonymous):

single term as in \[2x^{3}y\] not \[2x^{3}y-4xy^{2}+9x^{3}y^{2}\]

OpenStudy (anonymous):

right? and the degree is 7

OpenStudy (anonymous):

Right and right.

OpenStudy (anonymous):

so the degre of the one i posted is.. 5?

OpenStudy (anonymous):

Yep, on the last term.

OpenStudy (anonymous):

cool! one more, if you can? Which binomial is a factor of \[6x ^{2}+x+12\]? 3x-3,3x-4, 6x+3, or x+4

OpenStudy (anonymous):

What do you wish to find?

OpenStudy (anonymous):

Idk the question is Which binomail is a facot of that thing that i posted up ther^

OpenStudy (anonymous):

You'll have to use the quadratic formula to factor this, unless you mean 6x^2+xMINUS 12, which I think you do.

OpenStudy (anonymous):

i do yteah, soorry

OpenStudy (anonymous):

Factor by decomposition. Find two numbers that multiply to 6*12 and add up to the coefficient of x, 1.

OpenStudy (anonymous):

when i factored it, i got (3x+4)(2x-3). thats not any of theanswers

OpenStudy (anonymous):

\[6x^2+x-12\] \[6x^2+9x-8x-12\] \[3x(2x+3)-4(2x+3)\] \[(3x-4)(2x+3)\]

OpenStudy (anonymous):

yeah i saw what i did wrong thanks ! :)